# Electroweak Monopole-Antimonopole Pair Production at LHC

### v4

@@ Abstract
The monopole production at LHC crucially depends on the monopole production mechanism. We show that, if the monopole production mechanism at LHC is the thermal fluctuation of the Higgs vacuum as the early universe did, it is practically impossible for LHC to produce the monopole, even with the future FCC energy upgrade. This is because the temperature of the p-p fireball is simply too low to produce the monopole. But if the monopole production mechanism is the Drell-Yan and/or Schwinger mechanism, the 14 TeV LHC could produce the monopole in the form of the monopolium of mass around 5,7 TeV even when the monopole mass becomes 11 TeV. Or, it could produce the monopole when the mass is less than 7 TeV. Our result tells that, if the monopole production mechanism at LHC becomes the thermal fluctuation, searching for the remnant monopoles produced in the early universe could be the only way to detect the electroweak monopole. We discuss the physical implications of our result.
electroweak monopole, monopole production mechanism at LHC, Drell-Yan process, Schwiner mechanism, thermal fluctuation of Higgs vacuum, Ginzburg temperature, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
@@ I Introduction
With the advent of the Dirac’s monopole the magnetic monopole has become an obsession in physics, experimentally as well as theoretically [1, 2]. After the Dirac monopole we have had the Wu-Yang monopole, the ’t Hooft-Polyakov monopole, and the grand unification monopole [3, 4, 5]. But the electroweak (Cho-Maison) monopole, stands out as the most realistic monopole that could exist in nature and could actually be detected [6, 7].
@@
This makes the experimental confirmation of the electroweak monopole one of the most urgent issues in the standard model after the discovery of the Higgs particle. In fact the detection of this monopole, not the Higgs particle, should be regarded as the final (and topological) test of the standard model. For this reason the newest MoEDAL detector at LHC is actively searching for the monopole [8, 9, 10, 11].
To detect the electroweak monopole at LHC, we need to recall basic facts about the monopole [6, 7, 12, 13, 14, 15, 16]. First, this is the monopole which exists within (not beyond) the standard model as the electroweak generalization of the Dirac monopole, which can be viewed as a hybrid between Dirac and ’t Hooft-Polyakov monopoles. Second, the magnetic charge of the monopole is not $2\pi/e$ but $4\pi/e$, twice that of the Dirac monopole. This is because the period of the electromagnetic U(1) subgroup of the standard model becomes $4\pi$. Third, the mass of the monopole $M$ is of the order of 10 TeV. This is because the mass basically comes from the same Higgs mechanism which makes the W boson massive, except that here the magnetic coupling makes the monopole mass $1/\alpha$ times heavier than the W boson mass, around 11.0 TeV [17]. Despite this, the size of the monopole is set by the W boson mass, because the monopole solution has the Higgs and W boson dressing which fixes the size by the W boson mass.
Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it. However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV. This is problematic, because the monopole mass could turn out to be bigger than this. If this is so, we may have little chance to detect the monopole at LHC, and may have to try to detect the remnant monopoles at present universe produced in the early universe [15, 18].
However, two observations could make the monopole production at LHC possible even when the mass becomes over 7 TeV. First, the monopole has to be created in pairs at LHC. This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole. In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass becomes bigger than 7 TeV.
@@
A unique feature of this topological thermal fluctuation mechanism is that in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [15, 18]. In this case there is a possibility that LHC could produce the electroweak monopole even when the zero temperature mass becomes heavier than 7 TeV.
The purpose of this paper is to discuss how LHC can accommodate the above ideas and produce the electroweak monopoles even when the monopole mass exceeds 7 TeV. We show that, if the monopole production mechanism at LHC is the Drell-Yan process (and/or Schwinger mechanism), the present 14 TeV LHC could produce the monopole if the mass is less than 7 TeV. Or, it could produce the monopole in the form of the monopolium bound states of mass around 5.7 TeV even when the monopole mass becomes 11 TeV. This is because the binding energy of the monopoliun could reduce the monopolium mass by 16.3 TeV. Indeed we argue that the LHC can produce the monopolium for any reasonable monopole mass, as far as the mass does not exceed 29.8 TeV. However, if the monopole production mechanism at LHC becomes the thermal fluctuation of the Higgs vacuum, it has practically no chance to produce the monopole even with the future FCC energy upgrade. This is because the temperature of the p-p fireball at LHC simply becomes too low to produce the monopole thermally. This makes the monopolium bound state a most probable signal for the monopole production at LHC.
The paper is organized as follows. In Section II we discussthe electroweak monopole production mechanism at LHC. In Section III we review the monopole production in the electroweak phase transition in the early universe. In Section IV we discuss if LHC could produce the monopole in a similar manner, and argue that this is practically impossible because the temperature of the p-p fireball is too low. In Section V we discuss the naive Bohr model of the electroweak monopolium to see whether LHC could produce such electroweak monopolium. In Section VI we discuss a more realistic electroweak monopolium that could be produced at LHC. In Section VII we discuss the physical impications of our results.
II Electroweak Monopole Production Mechanism at LHC
@@
Notice that here we have defined $A_{\mu}^{(em)}$ and $Z_{\mu}$ gauge independently, without choosing the unitary gauge.
@@
The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field. It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [6, 7]. Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV. Intuitively we could argue the mass to be $1/\alpha$ times bigger than the W boson mass, around 11 TeV [17]. This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$). We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [12, 13, 14, 15, 16]. But in the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
@@
To understand the electroweak monopole production at LHC we have to know the monopole production mechanism at LHC. In the Drell-Yan process the monopole-antimonopole are produced in pairs by the electromagnetic interaction through the photon [8, 9]. But in the Schwinger mechanism a strong magnetic field is supposed to create the monopole-antimonopole pair, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [10, 19]. The justification for this comes from the asumption that a theory of monopole must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to $(g,-e)$. The Drell-Yan process is shown graphically in Fig. 1.
@@
But in the topological monopole production, the thermal fluctuation of the Higgs vacuum induces the change of topology and produce the monopole, which is exactly the electroweak monopole production mechanism in the early universe [15]. In this picture the monopoles are produced after the phase transition and stops at the Ginzburg temperature [15]. The topological monopole production mecanism is shown graphically in Fig. 2 for comparison.
@@
Notice that, in the Drell-Yan process the monopole pair is produced by the photon solely by the electromagnetic process. But the Weinberg-Salam Lagrangian perturbatively has no interaction which can be described by the above Feynman diagram. Similarly, for the Schwinger mechanism the one-loop effective action of the Weinberg-Salam theory has no indication of the monopole pair production in strong magnetic background. Most importantly, we need the change of topology to produce the monopole, but this topological nature of the monopole has no role at all in these mechanisms.
@@
In comparison the topology plays an essential role in the thermal production of the monopole, because here the thermal fluctuation of the Higgs vacuum is precisely what we need to induce the change of topology. As a result the monopole production in this mechanism is not controlled by any fundamental interaction or fundamental constant. Perhaps more importantly, in this thermal fluctuation the monopole mass depends on the temperature at creation, the monopole mass at creation could be considerably smaller than the monopole mass at zero temperature. In this case we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass. So, from now on we call the initial monopole mass (the monopole mass at birth) as the “baby” (or “infant”) monopole mass and the zero temperature mass as the “adolescent” monopole mass.
@@
It should be emphasized that the above monopole production mechanisms at LHC are the logical possibilities, not the realistic possibilities. For example, the Schwinger mechanism is certainly a logical possibility. However, it is not clear at all if (and how) the p-p fireball at LHC could create a strong magnetic background which can actually produce the monopole-antimonopole pair.
@@
If so, what is the the monopole production mechanism at LHC? We do not know yet. On the other hand, it has often been claimed that LHC could reproduce the early universe. If this is true, the monopole production mechanism at LHC could be topological, just like the electroweak monopole production in the early universe. In the following we discuss if LHC could produce the electroweak monopole thermally.
III Monopole Production in Early Universe: A Review
@@
It has generally been believed that the monopole production in the early universe critically depends on the type of the phase transition. In the first order phase transition the vacuum bubble collisions in the unstable vacuum are supposed to create the monopoles through the quantum tunneling to the stable vacuum during the phase transition, so that the monopole production is supposed to be suppressed exponentially by the vacuum tunneling [21]. On the other hand the monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism with no exponential suppression [22, 23].
@@
To amplify this point we start from the temperature dependent effective action of the standard model which describes the electroweak monopole production in the early universe [15, 25, 26, 18]
@@
Figure 4: The determination of the Ginzburg temperature $T_{G}$ in the electroweak phase transition. Here the red and blue curve represents $\xi^{3}\Delta F$ and the black line represents $T_{G}$.
@@
To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [kolb]
@@
=$=\left\{\begin{array}[]{ll}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~{}\dfrac{M_{H}^{2}}{2},&~{}~{}~{}T>T_{c},\\ \Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~{}\dfrac{M_{H}^{2}}{2},&~{}~{}~{}T\leq T_{c},\end{array}\right.$ (35)
@@
This tells that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature. But the W-boson which is massless above $T_{c}$ (before the symmetry breaking) becomes massive at $T_{c}$, and we have
@@
This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$). This confirms that the baby monopole mass near $T_{c}$ could be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value. This could allow LHC to produce the monopole even when the adolescent mass is bigger than 7 TeV. In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
@@
IV Electroweak Monopole Production from p-p Fireball at LHC
If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it has been asserted, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does. In this case, the above discussion suggests that LHC could produce the baby electroweak monopole of mass around 1.0 TeV just after $T_{c}$ and produce the last monopole of mass 10.4 TeV at the Ginzburg temperature $T_{G}$.
However, the 14 TeV LHC can not produce the monopole of mass 10.4 TeV. This means that the present LHC can not produce the monopole exactly as the early universe does. In other words, although LHC could reproduce the early universe, it could do so only in a limited sense. This is because, unlike the big bang fueled by the infinite energy, LHC can provide only a finite energy. So we have to take care of this energy constraint at LHC.

### v5

@@ Abstract
It is argued that the monopole production at LHC crucially depends on the monopole production mechanism. We show that, if the monopole production mechanism at LHC becomesthe thermal fluctuation of the Higgs vacuum as the early universe did, it is practically impossible for LHC to produce the monopole. This is because the temperature of the p-p fireball is simply too low to generate the thermal fluctuation necessary for the monopole production. But if the monopole production mechanism becomes the Drell-Yan and/or Schwinger mechanism, the 14 TeV LHC could produce the monopole only when the mass is less than 7 TeV. To circumbent this energy constraint, we propose a new monopole production mechanism at LHC, the monopole production in the form of the monopolium bound state. We show that LHC could produce the monopole in the form of the monopolium of mass less than 14 TeV, even when the monopole mass becomes considerably larger than 11 TeV. In this case we can prove the existence of the electroweak monopole at present LHC by detecting the decay modes of the monopolium. This implies that the monopole production in the form of the monopolium could be the most probable way to detect the electroweak monopole at LHC. We discuss the physical implications of our result.
electroweak monopole, monopole production mechanism at LHC, Drell-Yan process, Schwiner mechanism, monopolre production by thermal fluctuation of Higgs vacuum, monopokle production in the form of the monopolium, Ginzburg temperature, baby monopole mass, adolescent monopole mass, electroweak monopolium, Bohr radius of electroweak monopolium, mass of the electroweak monopolium, Rydburg energy of electroweak monopolium, monopolium production at LHC
@@ I Introduction
With the advent of the Dirac’s monopole the magnetic monopole has become an obsession in physics, experimentally as well as theoretically [1, 2]. After the Dirac monopole we have had the Wu-Yang monopole, the ’tHooft-Polyakov monopole, and the grand unification monopole [3, 4, 5]. But the electroweak (“Cho-Maison”) monopole stands out as the most realistic monopole that could exist in nature and could actually be detected [6, 7].
@@
This makes the experimental confirmation of the electroweak monopole one of the most urgent issues in the standard model after the discovery of the Higgs particle [8, 9, 10]. In fact the detection of this monopole, not the Higgs particle, should be regarded as the final (and topological) test of the standard model. For this reason the MoEDAL and ATLAS detectors at LHC are actively searching for the monopole [11, 12, 13, 14].
To detect the electroweak monopole at LHC, we need to remember the basic facts about the monopole [6, 7, 8, 15, 16, 17, 18]. First, this is the monopole which exists within (not beyond) the standard model as the electroweak generalization of the Dirac monopole, which can be viewed as a hybrid between Dirac and ’tHooft-Polyakov monopoles. Second, the magnetic charge of the monopole is not $2\pi/e$ but $4\pi/e$, twice that of the Dirac monopole. This is because the period of the electromagnetic U(1) subgroup of the standard model becomes $4\pi$. Third, the mass of the monopole $M$ is of the order of 10 TeV. This is because the mass basically comes from the same Higgs mechanism which makes the W boson massive, except that here the magnetic coupling makes the monopole mass $1/\alpha$ times heavier than the W boson mass, around 11.0 TeV [19]. Despite this, the size of the monopole is set by the W boson mass, because the monopole solution has the W (and Higgs) boson dressing which fixes the size by the W boson mass.
Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it. However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV. This is problematic, because the monopole mass could turn out to be bigger than this, around 11 TeV. If this is so, we may have little chance to detect the monopole at LHC, and may have to try to detect the remnant monopoles at present universe produced in the early universe [8, 20].
However, two observations could make the monopole production at LHC possible even when the mass becomes heavier than 7 TeV. First, the monopole has to be created in pairs at LHC. This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and anti-monopole pair. In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass becomes bigger than 7 TeV.
@@
A unique feature of this topological thermal fluctuation mechanism is that in this production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [8, 20]. In this case there is a possibility that LHC could produce the baby electroweak monopole even when the zero temperature mass becomes heavier than 7 TeV.
The purpose of this paper is to discuss how LHC can accommodate the above ideas and produce the electroweak monopoles even when the monopole mass exceeds 7 TeV. We confirm that, if the monopole production mechanism at LHC is the Drell-Yan process (and/or Schwinger mechanism), the present 14 TeV LHC could produce the monopole only if the mass is less than 7 TeV. However, if the monopole production mechanism at LHC becomes the thermal fluctuation of the Higgs vacuum, it has practically no chance to produce the monopole (even with the future FCC). This is because the temperature of the p-p fireball at LHC simply becomes too low to produce the monopole thermally. We propose a new monopole production mechanism at LHC, the monopole production in the form of the monopolium, which could allow LHC to produce the electroweak monopole even when the monopole mass becomes larger than 7 TeV. In specific, we show that LHC could produce the monopolium bound states of mass around 5.7 TeV for any reasonable monopole mass, as far as the mass does not exceed 29.8 TeV. This is because the binding energy of the monopoliun could reduce the monopolium mass by 16.3 TeV. This makes the monopolium bound state a most probable signal for the electroweak monopole production at LHC.
The paper is organized as follows. In Section II we discuss two electroweak monopole production mechanisms at LHC, Drell-Yan process and Schwinger mechanism. In Section III we review the monopole production in the electroweak phase transition in the early universe. In Section IV we discuss if LHC could produce the monopole in a similar manner, and argue that this is practically impossible because the temperature of the p-p fireball is too low. In Section V we discuss the naive Bohr model of thelectroweak monopolium and show that this model is unrealistic to describe an electroweak monopolium. In Section VI we discuss a more realistic electroweak monopole production mechanism, the monopole production via monopolium, and argue that such monopolium bound state could be produced at LHC. In Section VII we discuss the physical impications of our results.
II Electroweak Monopole Production Mechanism at LHC: Drell-Yan Process and Schwinger Mechanism
@@
Notice that here we have defined $A_{\mu}^{(em)}$ and $Z_{\mu}$ gauge independently, without any reference to the unitary gauge.
@@
The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $CP^{1}$ field. It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [6, 7]. Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV. Intuitively we could argue the mass to be $1/\alpha$ times bigger than the W boson mass, around 11 TeV. This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$) [19]. We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [15, 16, 17, 8, 18]. But in the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
@@
To understand the electroweak monopole production at LHC we have to know the monopole production mechanism at LHC. In the Drell-Yan process the monopole-antimonopole are produced in pairs by the electromagnetic interaction through the photon [11, 12]. But in the Schwinger mechanism a strong magnetic field is supposed to create the monopole-antimonopole pair, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [13, 21]. The justification for this comes from the asumption that a theory of monopole must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to (g,e)(g,-e). The Drell-Yan process is shown graphically in Fig. 1.
@@
But in the topological monopole production, the thermal fluctuation of the Higgs vacuum induces the change of topology and produce the monopole, which is exactly the monopole production mechanism in the early universe [8]. In this picture the monopoles are produced after the phase transition and stops at the Ginzburg temperature. The topological monopole production mecanism is shown graphically in Fig. 2 for comparison.
@@
Notice that, in the Drell-Yan process the monopole pair is produced via the photon solely by the electromagnetic process. But the Weinberg-Salam Lagrangian has no interaction which can be described by the above Feynman diagram perturbatively. Similarly, for the Schwinger mechanism the one-loop effective action of the Weinberg-Salam theory has no indication of the monopole pair production in strong magnetic background. So these monopole production mechanisms are possible beyond (not within) the standard model. More importantly, these mechanisms do not take into account the change of topology necessary to produce the monopole, so that the topological nature of the monopole is completely neglected in these mechanisms.
@@
In comparison the topology plays an essential role in the thermal production of the monopole, because here the thermal fluctuation of the Higgs vacuum is precisely what we need to induce the change of topology. As a result the monopole production in this mechanism is not controlled by any fundamental interaction or fundamental constant. Perhaps more importantly, in this thermal fluctuation the monopole mass depends on the temperature at creation because the vacuum value of the Higgs field which determines the monopole mass does. This means that the monopole mass at creation could be considerably smaller than the monopole mass at zero temperature. In this case we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass. So, from now on we call the initial monopole mass (the monopole mass at birth) as the “baby” (or “infant”) monopole mass and the zero temperature mass as the “adolescent” monopole mass.
@@
It should be emphasized that the above monopole production mechanisms at LHC are theoretical (i.e., logical) but not realistic possibilities. For example, the Schwinger mechanism is certainly a logical possibility. However, it is not clear at all if (and how) the p-p fireball at LHC could create a strong magnetic background which can actually produce the monopole-antimonopole pair.
@@
If so, what is the monopole production mechanism at LHC? We do not know yet. On the other hand, it has often been claimed that LHC could reproduce the early universe. If this is true, the monopole production mechanism at LHC could be thermal, just like the electroweak monopole production in the early universe. To discuss if this is true, we have to understand the monopole production mechanism in the early universe.
III Electroweak Monopole Production in the Early Universe: A Review
@@
It has generally been believed that the monopole production in the early universe critically depends on the type of the phase transition. In the first order phase transition the vacuum bubble collisions in the unstable vacuum are supposed to create the monopoles through the quantum tunneling to the stable vacuum during the phase transition, so that the monopole production is supposed to be suppressed exponentially by the vacuum tunneling [23]. On the other hand the monopole production in the second order phase transition is supposed to be described by the Kibble-Zurek mechanism which has no such exponential suppression [24, 25].
@@
To amplify this point we start from the temperature dependent effective action of the standard model which describes the electroweak phase transition and the monopole production in the early universe [8, 27, 28, 20]
@@
Figure 4: The determination of the Ginzburg temperature $T_{G}$ in the electroweak phase transition. Here the red and blue curve represents $\xi^{3}\Delta F$ and the temperature of the universe, and the black line represents $T_{G}$.
@@
To translate this in time scale, remember that the age of the universe $t$ in the radiation dominant era is given by [29]
@@
$=\left\{\begin{array}[]{ll}\Big{[}(\dfrac{T}{T_{2}})^{2}-1\Big{]}~\dfrac{M_{H}^{2}}{2},&~~~T>T_{c},\\ \Big{[}1+(\dfrac{\rho_{+}}{\rho_{0}})^{2}-(\dfrac{T}{T_{2}})^{2}\Big{]}~\dfrac{M_{H}^{2}}{2},&~~~T\leq T_{c},\end{array}\right$. (35)
@@
This confirms that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature. But the W-boson which is massless above $T_{c}$ (before the symmetry breaking) becomes massive at $T_{c}$, and we have
@@
This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~\text{TeV}$). This tells that the baby monopole mass near $T_{c}$ could be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value. This could allow LHC to produce the monopole even when the adolescent mass is bigger than 7 TeV. In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
$T$ $\rho_{+}(T)$ $\bar{M}_{H}(T)$ $\bar{M}_{W}(T)$ $\bar{M}(T)$
$T_{1}$ $146.74$ $16.3$ $5.9$ 0 0
$T_{c}$ $146.70$ $21.7$ $5.5$ $7.1$ $971.7$
$T_{2}$ $146.42$ $32.5$ $11.7$ $10.6$ 1 454.81\,454.8
$T_{i}^{LHC}$ $132.89$ $119.2$ $56.8$ $38.9$ 5 331.55\,331.5
$T_{min}^{LHC}$ $119.36$ $156.5$ $76.2$ $51.1$ 7 000.07\,000.0
$T_{i}$ $101.96$ $188.4$ $92.9$ $61.5$ 8 428.28\,428.2
$T_{G}$ $57.49$ $232.9$ $116.9$ $76.0$ 10 419.610\,419.6
0 $0.00$     $246.2$     $125.3$     $80.4$ 11 014.511\,014.5
@@
IV Topological Monopole Production by thermal fluctuation of Higgs vacuum at LHC
If the fireballs made of the p-p (and heavy ion) collisions at LHC can reproduce the hot thermal bath of radiation of the early universe as it has often been asserted, we may assume that LHC produces the electroweak monopoles by the same mechanism that the early universe does. In this case, the above discussion suggests that LHC could produce the baby electroweak monopole of mass around 1.0 TeV just after $T_{c}$ and produce the last monopole of mass 10.4 TeV at the Ginzburg temperature $T_{G}$.
However, the 14 TeV LHC can not produce the monopole of mass 10.4 TeV as it has to produce the monopole in pairs. This means that the present LHC can not produce the monopole exactly as the early universe does. In other words, although LHC could reproduce the early universe, it could do so only in a limited sense. This is because, unlike the big bang fueled by the infinite energy, LHC can provide only a finite energy. So we have to take care of this energy constraint at LHC.